OpenStudy (daniel.ohearn1):

Evaluate the integral from -inf to inf of ( 1 / (x^2 +1)^2 ) dx ?

OpenStudy (chillout):

First, do you know how to take the primitive?

OpenStudy (chillout):

Or rather, do you know improper integrals and simmetry?

OpenStudy (daniel.ohearn1):

Yes, improper integrals and symmetry.

OpenStudy (astrophysics):

Looks so arctany to me

OpenStudy (kainui):

I can show you how to solve this by an integral in the complex plane, but only if this is a complex analysis course. I think there's probably some way to do it with trig substitution though.

OpenStudy (astrophysics):

Yeah x = tanu should do the trick

OpenStudy (chillout):

You should get \(sec^{4}(u)\)

OpenStudy (chillout):

Actually as the function is even you should rewrite the integral first as \(\large2\int_{0}^{\infty}f(x)dx\)

OpenStudy (chillout):

Having trouble with the trig substitution?

OpenStudy (daniel.ohearn1):

is that (secxtanx )^2

OpenStudy (daniel.ohearn1):

Nom, no uh yea taking a bit

OpenStudy (chillout):

As you sunstitute \(x=tan(u)\) you get \(dx=sec^{2}(u)du\). Plugging in the substitution in the function we will have \((tan(u)^2+1\)). We recognize that the inside function as a fundamental trig relation, \(tan^{2}(u)+1=sec^2(u)\).

OpenStudy (chillout):

@Kainui, feel free to give your input, you're way better than me XD

OpenStudy (kainui):

Oh I am afk I have barely read the question it looks fine to me though!

OpenStudy (chillout):

Any advances so far?

OpenStudy (daniel.ohearn1):

so cos^2 x revaluate as a limit..

OpenStudy (chillout):

You're getting there!

OpenStudy (chillout):

I'll be afk for a bit.

OpenStudy (daniel.ohearn1):

Need I in Indeterminate form?

OpenStudy (daniel.ohearn1):

It blew up

OpenStudy (daniel.ohearn1):

Trading off into complex tinkering...

OpenStudy (astrophysics):

What does your integral look like as of right now?

OpenStudy (samigupta8):

Is the ans π/2?

OpenStudy (daniel.ohearn1):

gtg for now... later

OpenStudy (vishweshshrimali5):

hmmm... As @ChillOut aptly mentioned.. we will be using the substitution x = tan(u)...

OpenStudy (vishweshshrimali5):

That will give you dx = \(\sec^2 t \ dt\)...